The Feline Company

The Feline Company has been having some difficulties estimating its manufacturing overhead costs. In the past, manufacturing overhead costs have been related to production levels. However, some production managers have indicated that the size of their production lots might also be having an impact on the amount of their monthly manufacturing overhead costs. In order to investigate this possibility, the company collected information on its monthly manufacturing overhead costs, production in units, and average production lot size for 2020.
 

Month

Production   (Units)


Manufacturing   Overhead Cost


Average   Monthly Production Lot Size


1


75,000



$

925,800



20


2


90,000




843,875



19


3


65,000




910,125



24


4


80,000




946,000



19


5


55,000




879,000



24


6


50,000




825,000



18


7


85,000




960,000



22


8


105,000




1,053,500



25


9


102,000




1,020,000



23


10


68,000




905,000



20


11


75,000




938,000



22


12


95,000




995,000



24


 
 


Required:

(a.) Use the high-low method to estimate next month's manufacturing overhead costs, assuming the company is planning to produce 92,000 units


(b.) Use the high-low method to estimate next month's manufacturing overhead costs, assuming the company is planning to run a 21-lot size.
Answer #1

a)


CostActivity (Production in units)
Highest activity [A]1053500105000
Lowest activity [B]82500050000



Change in cost/activity [A-B]22850055000

Variable cost per unit = Change in cost /change in units

                         = 228500/55000

                          = $ 4.15455 per unit


Total cost-Variable cost (variable cost per unit * units )=Fixed cost
Highest activity1053500
4.15455*105000=436228
617272
Lowest activity825000
4.15455*50000= 207728
617272

At 92000Units ,Total manufacturing overhead cost = [4.15455*X] + 617272

Where X= number of units to be produced

                                                                  = [4.15455*92000 ] +617272

                                                                   = 382219 +617272

                                                                   = $ 999491

Total manufacturing cost to produce 92000 units = $ 999491

b)


CostActivity (Lot size)
Highest activity [A]105350025
Lowest activity [B]82500018



Change in cost/activity [A-B]2285007

Variable cost per unit = Change in cost /change in lot size

                     = 228500 /7

                     = $ 32642.85714 per lot size


Total cost-Variable cost (variable cost per unit * lot size )=Fixed cost
Highest activity1053500
32642.85714*25 = 816071
237429
Lowest activity825000
32642.85714 * 18= 587571
237429

At 21 Lot size ,Total manufacturing overhead cost = [32642.85714 * X] +237429

Where X= number of units to be produced

                                                                 = [32642.85714 * 21 ] + 237429

                                                                 = 685500+237429

                                                                  = $ 922929

Total manufacturing cost to produce 21 lot size = $ 922929


answered by: ANURANJAN SARSAM